You have $10,000 invested in a portfolio A. If you sell 70% and invest in an asset B whose return has a correlation of -0.5 with the return on portfolio. What is your overall portfolio variance then? Assume that the standard deviation of A is 7%, and that for B is 5%
Answer #1

Out of $10000, 7000(70%) is invested in asset B and the remaining i.e., 30% is invested in portfolio A

weight of portfolio A = wA = 0.3, standard deviation of A = σA = 7%

Weight of asset B = wB = 0.7, standard deviation of B = 5%+

Correlation between the returns of A and B = ρ = -0.5

Variance of the overall portfolio is calculated using the formula:

Variance of overall portfolio = σP2 = wA2A2 + wB2B2 + 2*ρ*wA*wBAB = 0.32*(7%)2 + 0.72*(5%)2 + 2*(-0.5)*0.3*0.7*7%*5% = 0.000441+0.001225+(-0.000735) = 0.000931

Answer -> Variance of overall portfolio = 0.000931